OpenStudy (anonymous):

Find the exact value... Sin20° Cos10° + Cos20° Sin10°

OpenStudy (anonymous):

Sin(20)Cos(10)+Cos(20)Sin(10) = sin(20+10) = sin 30 = 1/2

OpenStudy (anonymous):

@Kirby123 @siblings am i correct

OpenStudy (dayakar):

sinAcosB+cosAsinB = sin(A+B)

OpenStudy (dayakar):

here A=20,B=10

OpenStudy (anonymous):

I'm still confused

OpenStudy (anonymous):

@flowerfluff Correct!

OpenStudy (anonymous):

Oh never mind I get it now ok! Thanks!

OpenStudy (anonymous):

yw